OpenStudy (anonymous):

Write each of the following in terms of sin θ and cos θ, and then simplify if possible. csc θ tan θ

OpenStudy (anonymous):

\[\csc \theta = 1/\sin \theta\]

OpenStudy (anonymous):

\[\tan \theta = \sin \theta / \cos \theta\]

OpenStudy (anonymous):

the sin thetas cancel

OpenStudy (anonymous):

you get: \[1/\cos \theta = \sec \theta\]

OpenStudy (anonymous):

are there any other problems?

OpenStudy (anonymous):

sec θ tan θ csc θ this look hard

OpenStudy (anonymous):

you get: \[\frac{1}{\cos \theta} * \frac{\sin \theta}{\cos \theta} * \frac{1}{\sin \theta}\]

OpenStudy (anonymous):

the sin thetas cancel again

OpenStudy (anonymous):

you get: \[\frac{1}{\cos^2 \theta} = \sec^2 \theta\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome